What is the sum of the first 10 terms of the sequence defined by An=4n-3?
a. 37
b. 190
c. 380
d. 384

Answer by richwmiller(17219)

An=4n-3
A1=4-3=1
A10=4*10-3=37
S = ½(a + an)n
S = ½(1 + 37)10=190
b) 190